𝑠=𝑒𝑑+12π‘Žπ‘‘2 π‘’π‘ž(1) The tasks are to:

a) Plot a graph of distance (s) vs time (t) for the first 10s of motion if 𝑒=10π‘šπ‘ βˆ’1 and π‘Ž=5π‘šπ‘ βˆ’2.

b) Find the gradient of the graph at 𝑑=2𝑠 and 𝑑=6𝑠.

c) Differentiate the equation to find the functions for i) Velocity (𝑣=𝑑𝑠𝑑𝑑) ii) Acceleration (π‘Ž=𝑑𝑣𝑑𝑑=𝑑2𝑠𝑑𝑑2)


1
Expert's answer
2021-12-02T00:54:04-0500
"v(2)=20\frac{m}{s}\\v(6)=40\frac{m}{s}\\v(t)=10+5t\\a=5\frac{m}{s^2}"